OpenStudy (anonymous):

I don't know how to approach this question, please help! Find the value of the constant c for which the line y=4x+c is a tangent to the curve y^2=4x

OpenStudy (anonymous):

This is what I tried but it doesn't seem to be going anywhere.

OpenStudy (rock_mit182):

@ganeshie8 help us out

ganeshie8 (ganeshie8):

so far so good

ganeshie8 (ganeshie8):

\(16x^2-4x =-8xc-c^2 \) rearrange the equation in standard form \(16x^2+x(8c-4) + c^2 = 0\) For the line to be a tangent, you want the discriminant to be \(0\) : \((8c-4)^2 - 4*16*c^2=0\) solve \(c\)

OpenStudy (anonymous):

I am a bit confused with how you rearranged it.

OpenStudy (anonymous):

How did you decide to put those numbers in the brackets?

ganeshie8 (ganeshie8):

I wanted to see a quadratic in \(x\), so arranged it in decreasing exponent of \(x\)

ganeshie8 (ganeshie8):

Notice that the equation is in form \(ax^2+bx+c=0\)

OpenStudy (anonymous):

Ok, I think I understand better now. So if I started working on it from that last line I would be going in the right direction with it?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

yw